OpenStudy (anonymous):

Evaluate the limit and justify each step by indicating the appropriate properties of limites Lim 8x^2-9x+1/3x^2+9x-3 as x approaches Infinity @Mathematics

OpenStudy (anonymous):

nth term is the property limit = 8/3

OpenStudy (anonymous):

As long as the expression is (8x^2-9x+1)/(3x^2+9x-3) then I agree as with expressions that are written as P(x)/Q(x) where P(x) and Q(x) are polynomials then the limit as x->infinity is the leading terms of the numerator and denominator. In this case 8x^2/3x^2 = 8/3